3. A car headlight may put out 100 W at 25 V.
a) Calculate the current drawn by the headlight.
b) How many electrons pass through the bulb filament each second?
1
Expert's answer
2020-04-09T09:25:25-0400

a)


"I=\frac{P}{U}=\frac{100}{25}=4\ A"

b)


"N=\frac{I}{e}=\frac{4}{1.6\cdot 10^{-19}}=2.5\cdot 10^{19}"